OpenStudy (anonymous):

What does this mean? Evaluate sec 240° without using a calculator by using ratios in a reference triangle.

Parth (parthkohli):

First, \(\sec(240^{\circ}) = \dfrac{1}{\cos(240^\circ)}\). \(\cos(240^{\circ}) = \cos(180^{\circ}+ 60^{\circ}) = -\cos(60^{\circ})\)

Parth (parthkohli):

So we have boiled the problem down to\[\dfrac{1}{\cos(240^{\circ})} = \dfrac{1}{-\cos(60^{\circ})}\]

OpenStudy (anonymous):

Got you so far.

Parth (parthkohli):

What is \(\cos(60^{\circ})\)? If you need a reference triangle, here it is: